OpenStudy (anonymous):

Write the equation of the circunference with center C(-1, -5), and radius 7.

OpenStudy (anonymous):

Hi, I don't know exactly how to begin

OpenStudy (anonymous):

A circle is the set of points equidistant from a point C(h,k) called the center. The fixed distance r from the center to any point on the circle is called the radius. The standard equation of a circle with center C(h,k) and radius r is as follows: (x - h)2 + (y - k)2 = r2

OpenStudy (anonymous):

Clarifying that it is "squared": (x - h)^2 + (y - k)^2 = r^2

OpenStudy (anonymous):

Ok, but my h and k are negative

OpenStudy (anonymous):

Then you get -(-h) = +h.

OpenStudy (anonymous):

i get this: (x-1)^2+(y-5)^2=7^2

OpenStudy (anonymous):

It is +1 and +5 because (-(-1)) = +1 and (-(-5)) = +5

OpenStudy (anonymous):

ahhh i got it!

OpenStudy (anonymous):

Thank U!